OpenStudy (hellokitty17):

FAN AND MEDAL Factor completely 2x3 + 14x2 + 6x + 42.

oregonduck (oregonduck):

2[(x2 + 3)(x + 7)]

OpenStudy (hybrik):

To can factor this by grouping

OpenStudy (hellokitty17):

yup

OpenStudy (hybrik):

2x^2(x+7)+6(x+7) = (2x^2+6)(x+7)

OpenStudy (anonymous):

2x3=6 14x2=28+6x +42 6 + 28= 34 + 42= 76 76 + 6x (---- answer

OpenStudy (hybrik):

2x^2(x+7)+6(x+7) = (2x^2+6)(x+7) is correct.

OpenStudy (hybrik):

Wait

oregonduck (oregonduck):

no 2[(x2 + 3)(x + 7)] is correct

OpenStudy (hybrik):

You can simplify more

OpenStudy (hybrik):

2(x^2+3)(x+7)

OpenStudy (hybrik):

Oregon it is correct but usually you have to explain

OpenStudy (anonymous):

yes and then you can use foil

OpenStudy (hybrik):

To check it.

OpenStudy (hybrik):

Medal?

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

for team work